OpenStudy (anonymous):

If you people would be so kind as to answering a VERY important question for me. Z varies jointly with x and y. When x-2 and y-2, z-7 Please help me:/ I will give fan and medal... @Kristen17

OpenStudy (anonymous):

@wolfe8

OpenStudy (anonymous):

You mean when x = 2 and y = 2, z = 7, right? Z varies jointly with x and y means: Z = xyc for some constant c To figure out c plug in the x,y and z you know. 7 = 2*2*c you solve for c and get c = 7/4 so the equation is Z = 7/4 * xy

OpenStudy (anonymous):

Idk:/ That's the part that confused me. The paper is a little blurry but I think it has the - instead of an equal sign...

OpenStudy (anonymous):

Gotta be =. Makes no sense if it's a minus sign.

OpenStudy (anonymous):

Hmm okay. Thank you for looking at it:)

OpenStudy (anonymous):

Unless it's saying when x is -2 and y is -2, z is -7. In that case you get c = - 7/4 and your equation winds up being Z = - 7/4 * xy

OpenStudy (anonymous):

Nope. No "is" so guess it is an equal sign. Oh well. Thanks anyways